Consider a relation R= {(1, 1) (1, 3), (2, 2), (2,3) (3,123 on the set A = {1,2,37 Find transitive using warshalls algorithm... closure of the relation R consider

1
Expert's answer
2021-12-21T12:38:05-0500

the transitive closure of a binary relation R on a set X is the smallest relation on X that contains R and is transitive


 steps of Warshall’s Algorithm:

Step 1: Execute "W:=M_R, k:=0"

Step 2: Execute "k:=k+1"

Step 3: For all "i\ne k" such that "w_{ik}=1", and for all "j" execute the operation "w_{ij}:=w_{ij}\lor w_{kj}"

Step 4: If "k=n", then stop: the solution is "W=M_{R^*}", else go to step 2.


in our case:

"M_R=\begin{pmatrix}
 1 &0&1 \\
 0 & 1&1\\
1&0&0
\end{pmatrix}"


k = 1:

"w_{31}=1"


"W_1=\begin{pmatrix}
 1 &0&1 \\
 0 & 1&1\\
1&0&1
\end{pmatrix}"


k =3:

"w_{13}=1"


"W_2=\begin{pmatrix}
 1 &0&1 \\
 0 & 1&1\\
1&0&1
\end{pmatrix}"


"w_{23}=1"


"W_3=\begin{pmatrix}
 1 &0&1 \\
 1 & 1&1\\
1&0&1
\end{pmatrix}"


transitive closure:

"R^*=\{(1, 1) (1, 3), (2, 2), (2,3), (3,1),(2,1),(3,3)\}"